OpenStudy (anonymous):

e^(11 x -13) = 9 Find the exact value of x. No decimal entries. I got ln(9+13)/11 what did I do wrong?

OpenStudy (xishem):

\[e^{11x-13}=9\]\[\ln(e^{11x-13})=\ln(9)\]\[11x-13=\ln(9)\]Do you see now?

OpenStudy (anonymous):

You take the natural log of both sides, getting rid of the e, so now I solve from here?

OpenStudy (anonymous):

I still get the same answer ln(9)+13/11

OpenStudy (xishem):

\[\frac{\ln(9+13)}{11}\]Is the not the same as:\[\frac{\ln(9)+13}{11}\]

OpenStudy (agent0smith):

ln(9+13)/11 not the same answer.

OpenStudy (anonymous):

Now I see my mistake thanks guys